OpenStudy (taylor.joe):

Can someone help me with this problem? Paulina is remodeling her bathroom. The tile she has chosen is shown below. There are squares and trapezoids in the tile. The side length of each square in the tile is x centimeters. The height and the length of one of the bases of each trapezoid is x centimeters. The other length is 2x centimeters. Write a simplified equation to solve for x in terms of AT, the area of the tile. If necessary, use rational coefficients instead of root symbols.

OpenStudy (dani_rose):

Is there a pic of the tile?

OpenStudy (taylor.joe):

Yes.

OpenStudy (welshfella):

the area of each sqqqqquare = x^2 area of each trapezoid = (1/2) h (x + 2x) = (1/2) x*3x

OpenStudy (taylor.joe):

Thank you so much! Im new here on the site so i heard i need to give out medals or something? How do i do that? :)

OpenStudy (welshfella):

just click on the blue rectangle ' Best Response'

OpenStudy (welshfella):

Do you need anymore help?

OpenStudy (taylor.joe):

Oh! Okay! Thank you so much for your help.

OpenStudy (welshfella):

yw

OpenStudy (taylor.joe):

Oh well there is other questions to the problem but thanks to you i might actually figure them out. Thanks again! :)

OpenStudy (welshfella):

no probs